If n divided by m = 7/8, then what is m divided by n =. please show how u got the answer.
Answer by Alan3354(69443)

If n divided by m = 7/8, then what is m divided by n =. please show how u got the answer.
n/m = 7/8
Invert both sides
m/n = 8/7